OpenStudy (anonymous):

whats the probability of pulling a red card and a black card from a deck of cards?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

a 50% chance right?

OpenStudy (anonymous):

probably

OpenStudy (anonymous):

that seems to simple..

OpenStudy (kropot72):

\[P(red)=\frac{26}{52}\] \[P(black)=\frac{26}{51}\] \[P(red\ with\ black)=\frac{26\times 26}{52\times 51}\]